OpenStudy (anonymous):

2/1/3 x 3/1/2 , mixed number.

OpenStudy (bahrom7893):

36

OpenStudy (bahrom7893):

2/1/3=2*3/1=6

OpenStudy (bahrom7893):

3/1/2=3*2/1=6

OpenStudy (bahrom7893):

6x6=36

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

Is this 2 1/3 x 3 1/2?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

When I change it into improper fraction. it's like this. 7/3 x 7/2

OpenStudy (anonymous):

then I got stuck.

OpenStudy (anonymous):

ah okay so now you multiply 7x7 and 3x2 making it 49/6 which = 8 1/6 :)

OpenStudy (anonymous):

oh, okay. thanks.